# Risk-Sensitive Agent Compositions

### v2

@@ 3 Finding the path minimizes the value-at-risk
Towards optimizing the RMAG objective over an agent graph $G$, we remark that a simple procedure to find the optimal path is to estimate the $\textnormal{VaR}_{\alpha}$ along each path and pick the path that minimizes it. We can estimate the $\textnormal{VaR}_{\alpha}$ by computing the empirical $(1-\alpha)$-quantile on samples drawn from the loss along a path $L_{p}(Z_{p})$. While this is asymptotically optimal, it can be inefficient when the graph has an exponential number of paths as a function of the number of vertices.
@@
To understand how it works, we first present how we can estimate the value-at-risk of the loss distribution of a single path incrementally. To this end, consider a path in an agent graph $p=v_{1}\xrightarrow{e_{1}}\dots\xrightarrow{e_{m}}v_{m+1}$ with $v_{1}$ being the source vertex and $v_{m+1}$ the target vertex. Let $(Z_{1},\dots,Z_{m})\sim\mathcal{D}(T_{e_{1}}\times\cdots\times T_{e_{m}})$ be a sequence of random variables representing the sequences of traces of the agents along the path. Then define $R_{1},\dots,R_{m}$ to be real-valued random variables representing the losses incurred by the agents along each edge. Formally, we can write $R_{1}\coloneqq L_{e_{1}}(Z_{1}),\dots,R_{m}\coloneqq L_{e_{m}}(Z_{m})$. If $L_{p}$ is the composed loss function along the path as defined in Equation˜2, we can rewrite $L_{p}(Z)$ as $\max(R_{1},\dots,R_{m})$. A key observation that enables the incremental estimation of $\textnormal{VaR}_{\alpha}[L_{p}(Z)]$ (or equivalently, the $(1-\alpha)$-quantile of $L_{p}(Z)$) is that we can estimate $\textnormal{VaR}_{\alpha_{i}}[R_{i}]$ of each loss variable individually with $\alpha_{i}$ chosen such that $\sum\alpha_{i}=\alpha$. Following this, we can apply the union bound to obtain
@@
This bound is justified in the proof of Theorem˜1. This method allows us to estimate the value-at-risk of each edge loss variable independently and then combine the estimates at the end by computing their maximum. However, we need to optimize the allocation of the risk budgets $\alpha_{i}$ along each edge. Intuitively, allocating a higher risk budget to higher edge losses would result in a lower value-at-risk for that edge since it would correspond to a lower quantile.
BucketedVaR optimizes this by searching for risk budget allocations in a discretized set of buckets $B=\{0,\frac{\alpha}{d},2\frac{\alpha}{d},\dots,\alpha\}$ for a given discretization factor $d\in\mathbb{N}_{>0}$. At each vertex $v$ and corresponding to each bucket $\bar{\alpha}\in B$, the algorithm inductively builds an optimal path that minimizes the upper bound on the value-at-risk along the partial path at a risk budget of $\bar{\alpha}$. It achieves this by considering each predecessor $v^{\prime}$ of $v$ and each allocation of budget $\alpha^{\prime}\leq\bar{\alpha}$ until $v^{\prime}$, for which we have inductively constructed the optimal path, and then estimates the value-at-risk by allocating a budget of $\bar{\alpha}-\alpha^{\prime}$ to the loss along the edge $v^{\prime}\to v$. The optimal path for $(v,\bar{\alpha})$ is then obtained by extending the optimal path to $(v^{\prime},\alpha^{\prime})$ that minimizes the value-at-risk estimate. A detailed proof of soundness and run-time analysis of the algorithm can be found in the proof of Theorem˜1
@@
Observe that since $\gamma\to 0$ as $n\to\infty$, the theorem tells us that the estimated value-at-risk $q$ is at least as large as the true value-at-risk of the loss distribution of the path $p$, which corresponds to $\textnormal{quantile}(L_{p}(Z_{p}),1-\alpha)$. We can show this using a union bound argument for the budget allocations found by the algorithm along with the fact that the empirical CDF converges uniformly to the true CDF by the DKW inequality [8, 23]. The time complexity is easily deduced by looking at the pseudocode of BucketedVaR. See Appendix˜A for the complete proof. ∎
Theorem˜1 only justifies the validity of the union bounding method for estimating the value-at-risk but we have not discussed the optimality of BucketedVaR. Here, we can only hope to achieve tight estimates of the value-at-risk when the loss variables along every path are independent or are only loosely correlated. Under these independence assumptions, Theorem˜2 tells us that asymptotically, the quantile estimated by the BucketedVaR algorithm is at most the $(1-\alpha+\alpha^{2}/2)$-quantile of the losses along the optimal path. In other words, the path found by our algorithm is suboptimal by an additive factor of $\alpha^{2}/2$ with respect to the quantiles of the optimal path loss.
@@
This is again a consequence of the uniform convergence of the empirical CDF along with the fact that the independence of the loss variables allows us to deduce an upper bound. See Appendix˜A for the complete proof. ∎
@@ 4.1 Benchmarks
Maximizing safety during navigation and manipulation. We test our algorithm on the 16-Rooms and Fetch environments from Jothimurugan et al. [18] who present a compositional RL framework to complete long-horizon objectives in continuous environments. An objective is specified as a task graph which is the same as an agent graph. In the task graph, every edge represents a reach-avoid subtask: reaching a target region while avoiding a set of dangerous states. Each reach-avoid subtask specifies a real-valued reward function that takes in a trajectory of execution of the policy and evaluates to at least zero if the trajectory successfully completes the subtask. Intuitively, the reward captures a distance metric between trajectories and reach-avoid sets. The loss function in the agent graph is defined as the negative of the reward function.
The 16-Rooms environment consists of 16 rooms arranged in a $4\times 4$ grid with doors between adjacent rooms. RL policies are trained to control a point mass to complete reach-avoid tasks. The agent graph has 13 vertices and 16 paths of length 8 each. The second Fetch environment [26] is taken from Gymnasium-Robotics and involves controlling a 7-DoF robotic arm with a gripper attached at the end to pick up and move objects. We consider the task graph for the PickAndPlaceChoice objective that directs the gripper to move near the object, grip it, pick it up and move along one of two trajectories. The graph has 7 vertices and 2 paths of length 5. Additionally, we adapt the Rooms environment to implement a version of the drone navigation task from Example 1.
Figure 2: Top-view map with agent graph of the BoxRelay environment.
Minimizing resource consumption. We use the Miniworld [5] framework to implement the BoxRelay environment of which the top-view map is shown in Figure˜2. Observations in this environment are given as first-person view RGB images (see Section˜B.4). The objective is to move the player (red triangle) to the first box (marked 1) and pick it up. The player can then choose to go to either of the two boxes in the two middle rooms (3 or 7), place the initial box, pick up the next box, and finally drop it at the location marked in the right room with a star. Furthermore, after picking up the box at 7, the player can choose to exit through one of the two exits 8 or 9. We add the constraint that the player has limited battery power but gets recharged after dropping the first box and before picking up the second box. So we would like to minimize the amount of time that either box is held by the player. Importantly, since the initial positions of the player, the boxes, and the final target are all chosen uniformly at random within each room, this is a non-trivial optimization problem.
We can model the BoxRelay task as an agent graph as follows: the graph is as shown in Figure˜2 with the numbers being the vertices. The edges represent the agents which are RL policies trained to navigate the player to pick up and place objects at the desired locations. The loss functions evaluate to the number of time steps that a box is carried before being put back down and thus quantifies the resource consumption. Here, the RMAG objective corresponds to minimizing the maximum time that any box is carried by the player.
@@ 4.2 Experimental results
Equipped with these benchmark environments, we conduct three sets of experiments to evaluate the effectiveness of the approximation obtained from the BucketedVaR algorithm. We measure the accuracy of the approximation by estimating the empirical quantile on $10^{4}$ fresh samples of losses along the optimal path. Specifically, if $c\in\mathbb{R}$ is the approximate $\textnormal{VaR}_{\alpha}$ returned by the BucketedVaR algorithm, we draw samples of losses along the optimal path and compute the fraction of these which are below $c$ giving us the empirical quantile of $c$. We say that the approximation is tight if the empirical quantile is close to the desired quantile $1-\alpha$.
@@
Figure 3: Absolute quantile difference in percentage between the quantile computed by the BucketedVaR algorithm and the desired quantile. The empirical quantiles are computed on a fresh set of $10^{4}$ samples.
@@
Experiment 1: Comparison with optimal baseline algorithm. We define a baseline algorithm that explicitly esimates $\textnormal{VaR}_{\alpha}$ at the desired risk level $\alpha$ along each path of the agent graph. While this algorithm is asymptotically optimal, it is inefficient as discussed in the previous section. The number of samples for both algorithms is set to $10^{4}$ and the number of buckets is chosen based on the number of agents to allow sufficient granularity in the allocation of the risk budget. For example, in the 16-Rooms benchmark since there are 8 agents along each path, we choose a higher number of buckets.
@@
We find that BucketedVaR succeeds in finding the same optimal path as the baseline algorithm across all the benchmarks along with tight estimates of $\textnormal{VaR}_{\alpha}$. Figure˜3 shows the absolute difference in percentage points of the quantile estimates obtained by the BucketedVaR algorithm and the desired quantiles. We remark that the difference is never more than a couple percentage points across all benchmarks and risk levels. It also finds non-trivial allocations of the risk budget for the agents: in the 16-Rooms benchmark, the budget allocations for the 8 agents along the path to estimate $\textnormal{VaR}_{0.1}$ are $16\bar{\alpha},0\bar{\alpha},10\bar{\alpha},23\bar{\alpha},19\bar{\alpha},11\bar{\alpha},7\bar{\alpha}$, and $14\bar{\alpha}$ for $\bar{\alpha}=0.1/100$.
Experiment 2: Increasing sample size and number of buckets. Increasing the sample size improves the approximation because we have more accurate quantile estimates and similarly a higher number of buckets results in more granular risk budget allocations. This is confirmed by the two experiments we run: in the first one, plotted in Figure˜4(a), we maintain the number of buckets at 100 and vary the number of samples from 500 to $10^{4}$ with $\alpha=0.1$. The empirical quantile quickly stabilizes around 0.91 across benchmarks. The second experiment maintains the sample size at $10^{4}$ and varies the number of buckets from 5 to 100 with $\alpha=0.1$. Figure˜4(b) again shows a similar trend. We note that for the 16-Rooms benchmark, a low number of buckets gives bad VaR estimates since the number of agents along each path is higher.
Experiment 3: Scaling number of agents. From our theoretical result, we expect BucketedVaR to continue producing tight estimations of the VaR when we increase the number of agents along the path, provided that the condition on the independence of losses holds and the sample size is large enough. We test this hypothesis by simulating a long path of agents using the 16-Rooms benchmark. We consider the optimal path in this benchmark which has a sequence of 8 agents and construct a new agent graph with $8k$ agents for $k\in\{1,2,3,4,5\}$. We can sample a sequence of trajectories from this path by taking $k$ trajectories of the original 8 agents. The losses are then computed using the original loss functions. Since these are independently drawn trajectories, the assumption of independence holds. We estimate the $\textnormal{VaR}_{\alpha}$ of this path of agents using BucketedVaR for $10^{4}$ samples, $\alpha=0.1$, and $100$ buckets. The results are plotted in Figure˜4(c) in which we observe that the approximation continues to remain tight with an increasing number of agents.
Takeaways. Through our benchmarks, we illustrated that we can use our framework to model risk minimization associated with agent compositions with respect to safety and resource consumption requirements. Then we showed that the BucketedVaR algorithm produces estimates of the value-at-risk that are very close to the desired level of $1-\alpha$. Furthermore, we confirmed our theoretical result that the approximation improves with growing sample size and granularity of the risk budget division given by the number of buckets.
@@ 6 Conclusion and future directions
In summary, we introduced a framework for choosing optimal agent compositions to minimize the value-at-risk of losses that encode violations of safety, fairness, and privacy requirements. Our dynamic programming approach efficiently allocates the risk budget across agents, supporting tight approximations under mild independence assumptions. Empirical results on compositional RL benchmarks confirm its effectiveness in identifying optimal paths and quantifying tail behaviors, paving the way for more reliable agentic workflows.
A direction for the future is to incorporate the BucketedVaR algorithm in existing agentic frameworks. It is also possible to speed up the algorithm through a GPU implementation since the computation of the intermediate VaR estimates is highly parallelizable. While we present an efficient algorithm to minimize the value-at-risk in agent graphs, it is also an interesting future direction to extend our work to the estimation of other popular risk measures like the expected shortfall (CVaR) and the class of coherent risk measures [2]. Yet another an interesting direction is to explore whether we can develop online multi-armed bandit style algorithms to minimize risk. Such an algorithm would have to pick a composition of agents at each turn and minimize long-term regret with respect to the optimal composition. Recent works in risk-aware bandits [28] and combinatorial bandits [3] seem to indicate that this is feasible.
Appendix A Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
Proof of Theorem˜1.

### v3

@@ 3 Finding the path minimizing risk
We first consider the RMAG objective with the VaR risk measure and then show we can recover the CVaR through our procedure. A simple procedure to find the optimal path is to estimate the $\textnormal{VaR}_{\alpha}$ along each path and pick the path that minimizes it. We can estimate the $\textnormal{VaR}_{\alpha}$ by computing the empirical $(1-\alpha)$-quantile on samples drawn from the loss along a path $L_{p}(Z_{p})$. While this is asymptotically optimal, it can be inefficient when the graph has an exponential number of paths as a function of the number of vertices.
@@
To understand how it works, we first present how we can estimate the value-at-risk of the loss distribution of a single path incrementally. To this end, consider a path in an agent graph $p=v_{1}\xrightarrow{e_{1}}\dots\xrightarrow{e_{m}}v_{m+1}$ with $v_{1}$ being the source vertex and $v_{m+1}$ the target vertex. Let $(Z_{1},\dots,Z_{m})\sim\mathcal{D}(T_{e_{1}}\times\cdots\times T_{e_{m}})$ be a sequence of random variables representing the sequences of traces of the agents along the path. Then define $R_{1},\dots,R_{m}$ to be real-valued random variables representing the losses incurred by the agents along each edge. Formally, we can write $R_{1}\coloneqq L_{e_{1}}(Z_{1}),\dots,R_{m}\coloneqq L_{e_{m}}(Z_{m})$. If $L_{p}$ is the composed loss function along the path as defined in Equation 3, we can rewrite $L_{p}(Z)$ as $\max(R_{1},\dots,R_{m})$. A key observation that enables the incremental estimation of $\textnormal{VaR}_{\alpha}[L_{p}(Z)]$ (or equivalently, the $(1-\alpha)$-quantile of $L_{p}(Z)$) is that we can estimate $\textnormal{VaR}_{\alpha_{i}}[R_{i}]$ of each loss variable individually with $\alpha_{i}$ chosen such that $\sum\alpha_{i}=\alpha$. Following this, we can apply the union bound to obtain
@@
This bound is justified in the proof of Theorem 1. This method allows us to estimate the value-at-risk of each edge loss variable independently and then combine the estimates at the end by computing their maximum. However, we need to optimize the allocation of the risk budgets $\alpha_{i}$ along each edge. Intuitively, allocating a higher risk budget to higher edge losses would result in a lower value-at-risk for that edge since it would correspond to a lower quantile.
BucketedVaR optimizes this by searching for risk budget allocations in a discretized set of buckets $B=\{0,\frac{\alpha}{d},2\frac{\alpha}{d},\dots,\alpha\}$ for a given discretization factor $d\in\mathbb{N}_{>0}$. At each vertex $v$ and corresponding to each bucket $\bar{\alpha}\in B$, the algorithm inductively builds an optimal path that minimizes the upper bound on the value-at-risk along the partial path at a risk budget of $\bar{\alpha}$. It achieves this by considering each predecessor $v^{\prime}$ of $v$ and each allocation of budget $\alpha^{\prime}\leq\bar{\alpha}$ until $v^{\prime}$, for which we have inductively constructed the optimal path, and then estimates the value-at-risk by allocating a budget of $\bar{\alpha}-\alpha^{\prime}$ to the loss along the edge $v^{\prime}\to v$. The optimal path for $(v,\bar{\alpha})$ is then obtained by extending the optimal path to $(v^{\prime},\alpha^{\prime})$ that minimizes the value-at-risk estimate. A detailed proof of soundness and run-time analysis of the algorithm can be found in the proof of Theorem 1
@@
Observe that since $\gamma\to 0$ as $n\to\infty$, the theorem tells us that the estimated value-at-risk $q$ is at least as large as the true value-at-risk of the loss distribution of the path $p$, which corresponds to $\textnormal{quantile}(L_{p}(Z_{p}),1-\alpha)$. We can show this using a union bound argument for the budget allocations found by the algorithm along with the fact that the empirical CDF converges uniformly to the true CDF by the DKW inequality [Dvoretzky et al., 1956, Massart, 1990]. The time complexity is easily deduced by looking at the pseudocode of BucketedVaR. See Appendix B for the complete proof. ∎
Theorem 1 only justifies the validity of the union bounding method for estimating the value-at-risk but we have not discussed the optimality of BucketedVaR. Here, we can only hope to achieve tight estimates of the value-at-risk when the loss variables along every path are independent or are only loosely correlated. Under these independence assumptions, Theorem 2 tells us that asymptotically, the quantile estimated by the BucketedVaR algorithm is at most the $(1-\alpha+\alpha^{2}/2)$-quantile of the losses along the optimal path. In other words, the path found by our algorithm is suboptimal by an additive factor of $\alpha^{2}/2$ with respect to the quantiles of the optimal path loss.
@@
This is again a consequence of the uniform convergence of the empirical CDF along with the fact that the independence of the loss variables allows us to deduce an upper bound. See Appendix B for the complete proof. ∎
Equipped with the BucketedVaR algorithm to approximate $\textnormal{VaR}_{\alpha}$ over the agent graph, we show how we can also approximate the conditional value-at-risk $\textnormal{CVaR}_{\alpha}$ using the quantities computed by the algorithm. From the expression for $\textnormal{CVaR}_{\alpha}$ in Equation 2, it is easy to see that we can approximate the integral as
$\textnormal{CVaR}_{\alpha}[L_{p}(Z)]\approx\frac{1}{d}\sum_{k=1}^{d}\textnormal{VaR}_{k\frac{\alpha}{d}}[L_{p}(Z)].$ (8)
But notice that the BucketedVaR algorithm has already optimized values of $\textnormal{VaR}_{k\frac{\alpha}{d}}[L_{p}(Z)]$ over the agent graph for $1\leq k\leq d$ and so it suffices to take the average of these values. Finally, we observe that increasing the sample size and the number of buckets also improves the approximation of $\textnormal{CVaR}_{\alpha}$.
@@ 4.1 Benchmarks
In the first set of benchmarks, the objective is to maximize safety during navigation and manipulation. We adapt the 16-Rooms and Fetch environments from Jothimurugan et al. [2021] in which the long-horizon objective is specified as a task graph. In a task graph, each edge represents a reach-avoid subtask: reaching a target region while avoiding a set of dangerous states. To adapt this into an agent graph, we define loss functions for each edge subtask as the negative of the cumulative reward function which captures the reach-avoid specification. Additionally, we adapt the 16-Rooms environment to implement a version of the drone navigation task from Example 1. Secondly, we implement the BoxRelay benchmark using the Miniworld [Chevalier-Boisvert et al., 2023] framework in which we seek to minimize resource consumption. The objective is to move the player around while picking up boxes and dropping them at different locations. We add the constraint that the player has limited battery power but gets recharged each time a box is dropped off at a target. Importantly, since the initial positions of the player, the boxes, and the final target are all stochastic, this is a non-trivial optimization problem. To model the BoxRelay task as an agent graph, we define loss functions that evaluate to the number of time steps that a box is carried before being put back down and thus quantifies resource consumption. Complete benchmark descriptions are included in Appendix C.
@@ 4.2 Experimental results
Equipped with these benchmark environments, we conduct three sets of experiments to evaluate the effectiveness of the approximation obtained from the BucketedVaR algorithm. We measure the accuracy of the approximation by estimating the empirical quantile on $10^{4}$ fresh samples of losses along the optimal path. Specifically, if $c\in\mathbb{R}$ is the approximate $\textnormal{VaR}_{\alpha}$ returned by the BucketedVaR algorithm, we draw samples of losses along the optimal path and compute the fraction of these which are below $c$ giving us the empirical quantile of $c$. We say that the approximation is tight if the empirical quantile is close to the desired quantile $1-\alpha$. We also use empirical quantiles to compare approximations of $\textnormal{CVaR}_{\alpha}$.
@@
Figure 2: (a) Absolute quantile difference in percentage between the quantile computed by the BucketedVaR algorithm and the desired quantile. (b) Absolute quantile difference in percentage between the approximate $\textnormal{CVaR}_{\alpha}$ and that estimated by the baseline. The empirical quantiles are computed on a fresh set of $10^{4}$ samples.
@@
Experiment 1: Comparison with optimal baseline algorithm. We define a baseline algorithm that explicitly esimates $\textnormal{VaR}_{\alpha}$ and $\textnormal{CVaR}_{\alpha}$ at the desired risk level $\alpha$ along each path of the agent graph. While this algorithm is asymptotically optimal, it is inefficient as discussed in the previous section. The number of samples for both algorithms is set to $10^{4}$ and the number of buckets is chosen based on the number of agents to allow sufficient granularity in the allocation of the risk budget. For example, in the 16-Rooms benchmark since there are 8 agents along each path, we choose a higher number of buckets.
@@
We find that BucketedVaR succeeds in finding the same optimal path as the baseline algorithm across all the benchmarks along with tight estimates of both $\textnormal{VaR}_{\alpha}$ and $\textnormal{CVaR}_{\alpha}$. Figure 2 shows the absolute difference in percentage points of the quantile estimates obtained by the BucketedVaR algorithm and the baseline quantiles. We remark that the difference is never more than a couple percentage points across all benchmarks and risk levels. It also finds non-trivial allocations of the risk budget for the agents: in the 16-Rooms benchmark, the budget allocations for the 8 agents along the path to estimate $\textnormal{VaR}_{0.1}$ are $16\bar{\alpha},0\bar{\alpha},10\bar{\alpha},23\bar{\alpha},19\bar{\alpha},11\bar{\alpha},7\bar{\alpha}$, and $14\bar{\alpha}$ for $\bar{\alpha}=0.1/100$.
@@
Experiment 2: Increasing sample size and number of buckets. Increasing the sample size improves the approximation because we have more accurate quantile estimates and similarly a higher number of buckets results in more granular risk budget allocations. This is confirmed by the two experiments we run: in the first one, plotted in Figure 3(a), we maintain the number of buckets at 100 and vary the number of samples from 500 to $10^{4}$ with $\alpha=0.1$. The empirical quantile quickly stabilizes around 0.91 across benchmarks. The second experiment maintains the sample size at $10^{4}$ and varies the number of buckets from 5 to 100 with $\alpha=0.1$. Figure 3(b) again shows a similar trend. We note that for the 16-Rooms benchmark, a low number of buckets gives bad VaR estimates since the number of agents along each path is higher.
Experiment 3: Scaling number of agents. From our theoretical result, we expect BucketedVaR to continue producing tight estimations of the VaR when we increase the number of agents along the path, provided that the condition on the independence of losses holds and the sample size is large enough. We test this hypothesis by simulating a long path of agents using the 16-Rooms benchmark. We consider the optimal path in this benchmark which has a sequence of 8 agents and construct a new agent graph with $8k$ agents for $k\in\{1,2,3,4,5\}$. We can sample a sequence of trajectories from this path by taking $k$ trajectories of the original 8 agents. The losses are then computed using the original loss functions. Since these are independently drawn trajectories, the assumption of independence holds. We estimate the $\textnormal{VaR}_{\alpha}$ of this path of agents using BucketedVaR for $10^{4}$ samples, $\alpha=0.1$, and $100$ buckets. The results are plotted in Figure 3(c) in which we observe that the approximation continues to remain tight with an increasing number of agents.
Figure 4: (a) Empirical $\textnormal{VaR}_{0.1}$ coverage produced by BucketedVaR algorithm with respect to increasing correlation and path length. (b) Absolute coverage error (in percent) between $\textnormal{CVaR}_{0.1}$ estimate produced by BucketedVaR and the baseline estimate. This is also plotted against increasing correlation and path length.
Experiment 4: Robustness to correlated losses. To test the quality of the union bound approximations made by the BucketedVaR algorithm when losses are not independent, we setup an artificial agent graph which consists of a path of varying length along with loss functions that sample from correlated Gaussian distributions (complete description in Section C.3). This allows us to study the tightness of the approximation with increasing correlation ($\rho$) and path length. The results are plotted in Figure 4. While the approximation breaks down when we have complete correlation ($\rho=1$), it is interesting to note that the algorithm continues to produce reasonable approximations upto a correlation level of $\rho=0.5$.
@@ 6 Conclusion and future directions
In summary, we introduced a framework for choosing optimal agent compositions to minimize the value-at-risk of losses that encode violations of safety, fairness, and privacy requirements. Our dynamic programming approach efficiently allocates the risk budget across agents, supporting tight approximations. Empirical results on compositional RL benchmarks confirm its effectiveness in identifying optimal paths and quantifying tail behaviors, paving the way for more reliable agentic workflows.
A direction for the future is to incorporate the BucketedVaR algorithm in existing agentic frameworks. It is also possible to speed up the algorithm through a GPU implementation since the computation of the intermediate VaR estimates is highly parallelizable. While we present an efficient algorithm to minimize the value-at-risk in agent graphs, it is also an interesting future direction to extend our work to the estimation of other popular risk measures that are included in the class of coherent risk measures [Artzner et al., 1999]. Yet another interesting direction is to explore whether we can develop online multi-armed bandit style algorithms to minimize risk. Such an algorithm would have to pick a composition of agents at each turn and minimize long-term regret with respect to the optimal composition. Recent works in risk-aware bandits [Tan et al., 2022] and combinatorial bandits [Ayyagari and Dukkipati, 2021] seem to indicate that this is feasible.
7 Reproducibility statement
The source code along with detailed instructions to reproduce results and plots have been included in the supplementary materials. Further benchmark and experimental details are also included in Appendix C.
Appendix A Use of large language models (LLMs)
LLMs were used as code assistants in writing and debugging parts of the codebase. They were also used to polish the writing.
Appendix B Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
Proof of Theorem 1.